# DustNET: enabling machine learning and AI models of dusty plasmas

### 23 Mar 2026 v2

@@ VII Summary
Dusty plasmas constitute a fundamental component of the universe, spanning laboratory systems, planetary environments, and astrophysical media. As observational and experimental capabilities continue to advance, the resulting growth in data volume, dimensionality, and complexity is outpacing traditional analysis methods. A central challenge is therefore the development of new frameworks that can systematically integrate heterogeneous datasets and extract physically meaningful insight. In this context, data-driven and AI-assisted methodologies are emerging as essential complements to conventional physics-based approaches.
DustNET addresses this challenge by providing a scalable, community-driven framework for constructing standardized, multi-modal datasets tailored to machine learning and AI applications in dusty plasmas. By integrating laboratory measurements, space mission data, astrophysical observations, and synthetic datasets, DustNET enables the training of models that are both scalable and physically interpretable. A key contribution of DustNET is the establishment of a common data and benchmarking infrastructure, which facilitates reproducibility, cross-regime generalization, and collaborative model development. Such capabilities are critical not only for advancing fundamental science but also for supporting emerging applications, including lunar exploration under the Artemis program and future Mars missions.
From a modeling perspective, dusty plasma research is undergoing a transition toward greater integration, automation, and predictive capability. Recent advances emphasize model reduction, multiphase and multiphysics coupling, and uncertainty-aware prediction to address the intrinsic complexity and scale disparity of these systems. Reduced-order approaches—including kinetic approximations, moment closures, and machine-learning surrogates—are increasingly used to replace computationally intensive particle-based simulations. At the same time, multiphysics frameworks are incorporating plasma dynamics together with dust charging, sheath formation, thermionic emission, material interactions, and neutral gas coupling, enabling more realistic and comprehensive simulations across laboratory, industrial, and astrophysical environments. Hybrid strategies that combine first-principles models with data-driven inference further improve predictive accuracy and robustness.
Machine learning and AI are also transforming experimental analysis workflows. A key insight is that dusty plasma data are inherently structured, enabling specialized architectures to outperform generic approaches. For example, U-Net–based models improve particle detection in dense and noisy images, self-organizing maps provide robust matching and clustering, and PointNet-type networks enable accurate structural classification of plasma crystals. These methods significantly enhance tracking accuracy and enable systematic investigation of dynamical phenomena such as dust-acoustic waves, electrorheological effects, and phase transitions. Improved particle tracking and structural analysis, in turn, open new pathways for studying the formation, evolution, and long-term behavior of complex plasma systems.
Beyond improved accuracy, AI-based methods enable real-time or near-real-time data processing, which is essential for resource-constrained and dynamic environments such as space-based experiments. The integration of machine learning with advanced diagnostics—including stereoscopic imaging, high-speed cameras, and multi-modal sensing—supports time-resolved, three-dimensional reconstruction of particle dynamics. More broadly, the combination of DustNET datasets, physics-informed machine learning, and edge AI systems points toward closed-loop experimental workflows, in which data acquisition, analysis, and model updating are performed autonomously.
Overall, these developments signal a paradigm shift toward data-centric, physics-informed, and computationally efficient approaches to dusty plasma research. DustNET, together with multi-modal foundation models such as DUST-MAP and emerging AI agents, provides a unifying framework that bridges plasma physics, materials science, astrophysics, and artificial intelligence. This integrated approach not only addresses long-standing open problems but also enables new applications and discovery pathways, positioning dusty plasma research as both a beneficiary and a contributor to next-generation AI and scientific modeling.
@@ Acknowledgements.
LANL work was performed under the auspices of the U.S. Department of Energy (DOE) by Triad National Security, LLC, operator of the Los Alamos National Laboratory under Contract No. 89233218CNA000001, including DoE Office of Fusion Energy Sciences, and LANL Laboratory Directed Research and Development (LDRD) Programs. This work led by Justus Liebig University and THM University of Applied Sciences is supported by the German Aerospace Agency (DLR). We thank the German Aerospace Society for providing powerful PCs in accordance with the funding decision, which facilitated the studies. This project was supported by the German Federal Ministry of Economic Affairs and Climate Action under contract No. 50WK2270B. The work at Donghua University was supported by the National Natural Science Foundation of China under Grant No. 12035003. The work at Soochow University was supported by the National Natural Science Foundation of China under Grant No. 12175159. University of Michigan work is supported by the National Science Foundation ECLIPSE program award No. 2206039, the U.S. Department of Energy, Office of Fusion Energy Sciences, Grant DE-SC0018058, and Department of Energy Contract No. 89233218CNA000001. This research used resources of the Magnetized Plasma Research Laboratory at Auburn University. ET and SCT are supported by US Department of Energy Grant No. DE-SC-0019176 and by National Science Foundation, Plasma Physics program via grant 2308948. The work by Mississippi State University is supported by the National Science Foundation, Plasma Physics program via grant 2308947. This work is also supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under award number DE-SC-0021030.

### 09 Apr 2026 v3

@@ VII Summary
Dusty plasmas pervade the universe, from laboratory experiments and industrial systems to planetary environments and astrophysical media, yet their behavior remains difficult to capture within a unified predictive framework. As observational and experimental capabilities expand, dusty plasma research is entering a data-rich era in which traditional modeling approaches alone are no longer sufficient. This transition calls for a new paradigm that integrates physical understanding with data-driven discovery, enabling systematic interpretation across scales, environments, and modalities.
DustNET represents a foundational step toward this paradigm. As a community-driven, multi-modal data ecosystem, DustNET aims at standardize, curate, and integrate heterogeneous datasets spanning laboratory measurements, space missions, astrophysical observations, and synthetic data. Beyond serving as a repository, DustNET establishes a common infrastructure for training, benchmarking, and validating machine learning models, thereby enabling reproducibility, cross-domain generalization, and collaborative progress. In this sense, DustNET plays a role analogous to ImageNet for computer vision and deep learning, but with an explicit emphasis on physics-informed learning and scientific interpretability.
DustNET forms the data backbone for a broader AI-enabled framework built around multi-modal foundation models and autonomous agents. Foundation models, such as DUST-MAP, aim to learn unified representations of dusty plasma systems by integrating images, time-series measurements, simulations, and scientific text within a common latent space. These models provide a “world model” of dusty plasmas, capable of capturing multi-scale, multiphysics behavior in a manner that is both data-driven and constrained by physical principles. Complementing this capability, AI agents can operate on top of such models to enable closed-loop experimentation, dynamically adjusting diagnostics, guiding simulations, and optimizing experimental conditions in real time.
From a modeling perspective, we anticipate a convergence of reduced-order physics models, multiphysics simulations, and machine learning surrogates into hybrid frameworks that are both computationally efficient and physically grounded. Advances in geometric learning, physics-informed loss functions, and uncertainty quantification will further enhance model robustness and interpretability. At the experimental level, integration with edge AI systems and advanced diagnostics will enable real-time, three-dimensional, and multi-modal characterization of dusty plasma dynamics, including particle tracking, structural evolution, and emergent phenomena.
Together, these developments form a pathway toward a unified, data-centric, and physics-informed framework for dusty plasma science. By bridging fundamental plasma physics with artificial intelligence, DustNET, DUST-MAP, and AI agents collectively provide a platform for addressing long-standing open problems while enabling new capabilities in space exploration, fusion research, and astrophysical discovery. More broadly, dusty plasmas offer a uniquely rich testbed for next-generation AI systems grounded in physical reality, positioning the field as both a beneficiary of and a contributor to the future of scientific machine learning and AI.
@@ Acknowledgements.
LANL work was performed under the auspices of the U.S. Department of Energy (DOE) by Triad National Security, LLC, operator of the Los Alamos National Laboratory under Contract No. 89233218CNA000001, including DoE Office of Fusion Energy Sciences, and LANL Laboratory Directed Research and Development (LDRD) Programs. The work led by Justus Liebig University and THM University of Applied Sciences is supported by the German Aerospace Agency (DLR). We thank the research group led Markus H. Thoma and M. Kretschmer at Justus Liebig University Giessen for providing the data used in this study and the German Aerospace Society for providing powerful PCs in accordance with the funding decision, which facilitated the studies. The project was supported by the German Federal Ministry of Economic Affairs and Climate Action under contract No. 50WK2270B. The work at Donghua University was supported by the National Natural Science Foundation of China under Grant No. 12035003. The work at Soochow University was supported by the National Natural Science Foundation of China under Grant No. 12175159. University of Michigan work is supported by the National Science Foundation ECLIPSE program award No. 2206039, the U.S. Department of Energy, Office of Fusion Energy Sciences, Grant DE-SC0018058, and Department of Energy Contract No. 89233218CNA000001. The work at the  University of Greifswald is supported by the German Aerospace Agency (DLR) under grant Nos 50WM2161 and 50WM2561. This research used resources of the Magnetized Plasma Research Laboratory at Auburn University. ET and SCT are supported by US Department of Energy Grant No. DE-SC-0019176 and by National Science Foundation, Plasma Physics program via grant 2308948. The work by Mississippi State University is supported by the National Science Foundation, Plasma Physics program via grant 2308947. This work is also supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under award number DE-SC0026203. Work at Iowa was supported by the U.S. Department of Energy grant DE-SC0025444 and National Science Foundation grant No. PHY-2510501.